OpenStudy (anonymous):

Can someone please help with a math question? I will medal best answer and ill fan you!!!

OpenStudy (anonymous):

What is the approximate volume of the cone? Use 3.14 for pi? the base radius is 10cm and the height is 9 cm...please help me find the answer!

Vocaloid (vocaloid):

volume of cone = (1/3)(pi)r^2*h

OpenStudy (anonymous):

k so the answer would be 94 cm right?

OpenStudy (anonymous):

no your decimal is in the wrong place

OpenStudy (anonymous):

\[\frac{1}{3}\times 9\times 10^2\times \pi\] think you forgot to square the \(10\)

OpenStudy (anonymous):

you can plug the numbers in here https://www.google.com/search?client=ubuntu&channel=fs&q=volume+of+a+cone&ie=utf-8&oe=utf-8 all i did was google "volume of a cone"

OpenStudy (anonymous):

oh, ok thanks!!! bye!